# FRQI Pairs method for image classification using Quantum Recurrent Neural Network

### v1

@@ I Introduction
Quantum image processing is a promising field in quantum computing. Starting with Venegas-Andraca and Bose’s [venegas2003storing] qubit lattice representation for quantum image encoding, the description of quantum images came right after this [latorre2005image]. Quantum states describe patterns for two main reasons: to improve classification efficiency [wiebe2015quantum] and to provide valuable models for traditional issues [tanaka2007quantum]. The Flexible Representation for Quantum Images (FRQI) was first proposed in [le2011flexible] and was further developed in [le2011flexible_2]. This research focuses on quantum image encoding and quantum machine learning classification methods applied to the MNIST data set.
The field of quantum machine learning is expanding rapidly, and new methods emerge, sometimes inspired by the classical machine learning methods or some developed purely for quantum computers. Some of the recent methods are inspired by traditional machine learning techniques, e.g., quantum state vector machine (QSVM) [delilbasic2021quantum, li2015experimental, rana2022comparative, rebentrost2014quantum], quantum $k$-nearest neighbors (QKNN) [dang2018image, ruan2017quantum, wang2019improved] and quantum nearest mean classifier (QNMC) [santucci2017quantum, sergioli2017quantum, sergioli2018quantum]. In the family of deep methods: variational quantum circuits (VQC) [potempa2022comparing, skolik2021layerwise, wang2022development] inspired by classical neural networks, quantum tensor networks (QTN) [grant2018hierarchical, huang2021variational, lazzarin2022multi], quantum convolutional neural networks (QCNN) [cong2019quantum, huang2022variational, hur2022quantum, shende2004minimal, zheng2023design], random quantum neural networks (RQNN) [konar2022random] and quantum recurrent neural networks (QRNN) [bausch2020recurrent, potempa2022recognition].
@@ II Subject
The prototyping results of this paper are based on the thesis [potempa2022recognition]. The work proposes a novel approach to the classification of quantum-encoded images using QRNN [bausch2020recurrent] with the input data encoded in a quantum way, using the FRQI [le2011flexible]. The use of quantum encoding allows for a futuristic assumption that the classification is performed on some universal quantum computer where the encoded data is stored in a quantum memory.
However, such devices are not available in the Noisy Intermediate Scale Quantum (NISQ) [preskill2018quantum] era, and the disposal of the data preprocessing steps required every time an image is loaded into the quantum system would significantly reduce the classical processing overhead.
@@ II-A MNIST Database
The Modified National Institute of Standards and Technology (MNIST) database contains handwritten digits stored in the form of $28\times 28$ pixel images with 02550...255 values representing the pixel intensity. A scaled-down sample image is shown in Fig. 1. The data set has been chosen due to its wide use for benchmark purposes, hence allowing for broad comparison with classical and quantum methods. The recent criticism of MNIST benchmarks for quantum machine learning algorithms [bowles2024better] suggests that further research is needed to test the approach on different datasets and from other perspectives.
@@ II-B Flexible Representation of Quantum Images
The FRQI method proposed in [le2011flexible] allows efficient storage of single channel images using only $\mathcal{O}(\lceil\log_{2}n\rceil)$ qubits, where $n$ is roughly the side length of the image. The method also allows the easy manipulation of the image properties, thus placing it and its descendants as a versatile image storing and manipulation method for a quantum computer. The method has since been improved to support multichannel images in further works [khan2019improved, li2018color, nasr2021efficient]. Since the MNIST database used for classification includes only grayscale images, the FRQI method for encoding has been used for simplicity.
@@ II-C Quantum Recurrent Neural Networks
[bausch2020recurrent] presents the first QRNN model capable of performing complex tasks such as sequence learning and digit classification. This QRNN utilizes an enhanced version of a quantum neuron presented in [cao2017quantum] with amplitude amplification to create a nonlinear activation function. The model was tested on various tasks, including memorization, sequence prediction, and MNIST digits dataset classification, demonstrating its capability to handle high-dimensional training data. The QRNN implemented using PyTorch programming library [paszke2019pytorch] shows results that have an impact on quantum machine learning, especially in managing long sequence data without the gradient vanishing problem typical in classical RNNs.
@@ III-A Examined Models
The thesis [potempa2022recognition] introduces a quantum-specific approach to combine images encoded in a quantum way with QRNN [bausch2020recurrent]. The novelty of the presented approach is the use of input images encoded by the FRQI method as described in Section II-B. The data set used for prototyping is the MNIST digits, scaled down to $8\times 8$ for most cases.
@@ III-B FRQI Pairs Model
For comparison, the number of cells in a direct implementation of a QRNN, where each cell takes one pixel of the same image, as in [bausch2020recurrent], requires an exponentially larger number of cells, compared to 3, i.e., $(2^{n})^{2}=2^{2n}$. An example diagram of the FRQI Pairs model for a $4\times 4$ image is shown on Fig. 2.
@@
Figure 2: Schematic diagram of the FRQI Pairs Model [potempa2022recognition] for $4\times 4$ image. The approach uses all combinations of $\{c,x,y\}$ qubits as inputs to QRNN cells, where $x$, $y$ correspond to the X and Y qubits for position encoding.
Figure 3: Aggregated results of the four champion FRQI Pairs models training [potempa2022recognition]: loss function - categorical cross-entropy (left), accuracy (center) and average confusion matrix for the test set (right).
@@ IV Comparative analysis
By fulfilling the conditions, it is possible to compare gathered publications with [potempa2022recognition]. The summary of this section is presented in Tab. I.
@@
[wang2022development] 2000+1000 images [^1] Training+test set sizes,
@@
[konar2022random] 4-class dataset (24k+4k1), noised, classical deep embedding into 128 values Amplitude encoding RQNN 4-class: 97.20%2
[bausch2020recurrent] Full dataset (60k+10k1),
@@
[potempa2022recognition] Full dataset (60k+10k1),
@@
The final results of [potempa2022recognition] show that the proposed model is able to grasp the principles of the underlying data distributions. Its test accuracy is 74.6%. However, not ideal, its accuracy at the level of magnitude allowing reasonable comparison to three of the presented works, namely [siemaszko2023rapid], [wang2022development] and [wei2022quantum].
@@
The work presented in [siemaszko2023rapid] implements a QRNN, which was tested against a binary classification of the MNIST digits $3$ and $6$. The results confirm the QRNN utility for the classification of handwritten digits. However, they suggest that the classical LSTM model with a similar number of parameters performs better, which might suggest that the QRNN model is inefficient in image classification.
@@
[wei2022quantum] represents a QCNN, which is a different family of models, yet the model was trained to classify all ten digits, and except for the 0-padding to $32\times 32$ size, no initial data transformation was performed. The numbers of trainable parameters of the model from [potempa2022recognition] and [wei2022quantum] are at a similar level: 716 and 379, respectively. For both models, the final accuracy is also similar: 74.6% and 74.3%, while the data used to train the final model have a higher resolution in the case of [wei2022quantum]. Both results suggest that the QRNN model might have a higher learning potential than what was achieved in [potempa2022recognition], and may achieve even better results for full-sized images. From 1 and 3 we find that the $28\times 28$ FRQI encoding uses 11 qubits, so the circuit will use 15 or more qubits, depending on the number of memory lanes, while the number of cells will increase to 25. The increase in the number of cells and qubits will result in a higher number of the model’s trainable parameters. However, it will also increase the model’s training capacity and potentially improve the results.
@@
The authors of [wang2022development] presented a VQDNN model, which resembles a densely connected network. They managed to heavily reduce the number of qubits required to operate the network, providing a solution that uses only ten qubits for 10-class problems. For the features extracted using the PCA, the authors experimented with amplitude and angle encoding, as well as their combination. The total number of parameters for the deepest model was 430. The depth of the model increases the model’s training capacity, and the best model achieved approximately 80% accuracy. The proposed model of the VQDNN network proved its potential for experimentation to replace the internals of the QRNN cell from [potempa2022recognition].
@@
The use of PCA as the feature extraction method should also be considered in the further development of FRQI-QRNN models, as its impact was shown in [bausch2020recurrent] and [wang2022development]. However, the method might not be directly applicable to the problem [potempa2022recognition] is trying to solve, i.e., the quantum ML model operating on already quantum-encoded data. Despite the fact that the PCA method can help reduce the dimensionality of feature spaces, thus reducing the model complexity and allowing for higher performance, with the same number of trainable parameters–one can imagine that the data stored in future quantum memory would be encoded and compressed using different means. The applicability of the quantum version of PCA [lloyd2014quantum, ostaszewski2015quantum] or other methods such as [li2021resonant] to the feature preprocessing phase should also be explored.
@@
Other works that use the amplitude encoding method for the classification of MNIST datasets are [konar2022random, zheng2023design]. They both present classifier types different from QRNN, but the authors managed to achieve high test accuracy on limited datasets. Random QNN [konar2022random] shows high robustness against noisy data. However, the image embedding is performed by a classical densely connected layer, which, similarly to PCA, offloads some part of the solution to the classical part.
@@
[zheng2023design] proposed the Approximation Quantum State Preparation (AQSP) method for image encoding, which uses a simulated quantum circuit to train the image representation and has a time complexity of only $\mathcal{O}(n)$. The authors also combine the AQSP method with the proposed QCNN framework. The design of a hybrid recurrent network with convolutional and pooling input layers might help make the FRQI-QRNN architecture less dependent on the input image size and reduce the complexity of the recurrent part.
@@ V Conclusions
The new FRQI Pairs architecture, presented in [potempa2022recognition], requires an exponentially lower number of recurrent cells compared to [bausch2020recurrent] ($n^{2}$ vs. $2^{2n}$ 3) which may lead to shorter execution times thus a lower chance of decoherence, and a higher computational efficiency during the inference phase if applied to real-world problems.
@@
Another important feat of the FRQI Pairs method is that it managed to train on the full MNIST dataset and has proven performance comparable to other methods like [ostaszewski2015quantum, wang2022development]. Many other presented works struggle to capture the idea of MNIST data base, which is to create a dataset for benchmarking machine learning models and ideas based on reliable and repeatable dataset. MNIST is meant to be taken as a whole, to enable direct comparison with numerous classical models.
Although areas were left to improve compared to [bausch2020recurrent], the method has the potential to be extended by various preprocessing routines [bausch2020recurrent, li2021resonant, lloyd2014quantum, ostaszewski2015quantum] and layer/model architectures [wang2022development, wei2022quantum, zheng2023design]. The approach of the other works to the problem suggests that the research from [potempa2022recognition] should also be extended to a binary image classification case for broader comparison possibilities.

### v2

@@ I Introduction
Quantum image processing is a promising field in quantum computing. Starting with Venegas-Andraca and Bose’s [37] qubit lattice representation for quantum image encoding, the description of quantum images came right after this [13]. Quantum states describe patterns for two main reasons: to improve classification efficiency [41] and to provide valuable models for traditional issues [36]. The Flexible Representation for Quantum Images (FRQI) was first proposed in [15] and was further developed in [16]. This research focuses on quantum image encoding and quantum machine learning classification methods applied to the MNIST data set.
The field of quantum machine learning is expanding rapidly, and new methods emerge, sometimes inspired by the classical machine learning methods or some developed purely for quantum computers. Some of the recent methods are inspired by traditional machine learning techniques, e.g., quantum state vector machine (QSVM) [6, 19, 27, 28], quantum $k$-nearest neighbors (QKNN) [5, 29, 39] and quantum nearest mean classifier (QNMC) [30, 31, 32]. In the family of deep methods: variational quantum circuits (VQC) [24, 35, 38] inspired by classical neural networks, quantum tensor networks (QTN) [7, 9, 14], quantum convolutional neural networks (QCNN) [4, 8, 10, 33, 42], random quantum neural networks (RQNN) [12] and quantum recurrent neural networks (QRNN) [1, 25].
@@ II Subject
The prototyping results of this paper are based on the thesis [25]. The work proposes a novel approach to the classification of quantum-encoded images using QRNN [1] with the input data encoded in a quantum way, using the FRQI [15]. The use of quantum encoding allows for a futuristic assumption that the classification is performed on some universal quantum computer where the encoded data is stored in a quantum memory.
However, such devices are not available in the Noisy Intermediate Scale Quantum (NISQ) [26] era, and the disposal of the data preprocessing steps required every time an image is loaded into the quantum system would significantly reduce the classical processing overhead.
@@ II-A MNIST Database
The Modified National Institute of Standards and Technology (MNIST) database contains handwritten digits stored in the form of $28\times 28$ pixel images with 02550...255 values representing the pixel intensity. A scaled-down sample image is shown in Fig. 1. The data set has been chosen due to its wide use for benchmark purposes, hence allowing for broad comparison with classical and quantum methods. The recent criticism of MNIST benchmarks for quantum machine learning algorithms [2] suggests that further research is needed to test the approach on different datasets and from other perspectives.
@@ II-B Flexible Representation of Quantum Images
The FRQI method proposed in [15] allows efficient storage of single channel images using only $\mathcal{O}(\lceil\log_{2}n\rceil)$ qubits, where $n$ is roughly the side length of the image. The method also allows the easy manipulation of the image properties, thus placing it and its descendants as a versatile image storing and manipulation method for a quantum computer. The method has since been improved to support multichannel images in further works [11, 17, 21]. Since the MNIST database used for classification includes only grayscale images, the FRQI method for encoding has been used for simplicity.
@@ II-C Quantum Recurrent Neural Networks
[1] presents the first QRNN model capable of performing complex tasks such as sequence learning and digit classification. This QRNN utilizes an enhanced version of a quantum neuron presented in [3] with amplitude amplification to create a nonlinear activation function. The model was tested on various tasks, including memorization, sequence prediction, and MNIST digits dataset classification, demonstrating its capability to handle high-dimensional training data. The QRNN implemented using PyTorch programming library [23] shows results that have an impact on quantum machine learning, especially in managing long sequence data without the gradient vanishing problem typical in classical RNNs.
@@ III-A Examined Models
The thesis [25] introduces a quantum-specific approach to combine images encoded in a quantum way with QRNN [1]. The novelty of the presented approach is the use of input images encoded by the FRQI method as described in Section II-B. The data set used for prototyping is the MNIST digits, scaled down to $8\times 8$ for most cases.
@@ III-B FRQI Pairs Model
For comparison, the number of cells in a direct implementation of a QRNN, where each cell takes one pixel of the same image, as in [1], requires an exponentially larger number of cells, compared to 3, i.e., $(2^{n})^{2}=2^{2n}$. An example diagram of the FRQI Pairs model for a $4\times 4$ image is shown on Fig. 2.
@@
Figure 2: Schematic diagram of the FRQI Pairs Model [25] for $4\times 4$ image. The approach uses all combinations of $\{c,x,y\}$ qubits as inputs to QRNN cells, where $x$, $y$ correspond to the X and Y qubits for position encoding.
Figure 3: Aggregated results of the four champion FRQI Pairs models training [25]: loss function - categorical cross-entropy (left), accuracy (center) and average confusion matrix for the test set (right).
@@ IV Comparative analysis
By fulfilling the conditions, it is possible to compare gathered publications with [25]. The summary of this section is presented in Tab. I.
@@
[38] 2000+1000 images [^1] Training+test set sizes,
@@
[12] 4-class dataset (24k+4k1), noised, classical deep embedding into 128 values Amplitude encoding RQNN 4-class: 97.20%2
[1] Full dataset (60k+10k1),
@@
[25] Full dataset (60k+10k1),
@@
The final results of [25] show that the proposed model is able to grasp the principles of the underlying data distributions. Its test accuracy is 74.6%. However, not ideal, its accuracy at the level of magnitude allowing reasonable comparison to three of the presented works, namely [34], [38] and [40].
@@
The work presented in [34] implements a QRNN, which was tested against a binary classification of the MNIST digits $3$ and $6$. The results confirm the QRNN utility for the classification of handwritten digits. However, they suggest that the classical LSTM model with a similar number of parameters performs better, which might suggest that the QRNN model is inefficient in image classification.
@@
[40] represents a QCNN, which is a different family of models, yet the model was trained to classify all ten digits, and except for the 0-padding to $32\times 32$ size, no initial data transformation was performed. The numbers of trainable parameters of the model from [25] and [40] are at a similar level: 716 and 379, respectively. For both models, the final accuracy is also similar: 74.6% and 74.3%, while the data used to train the final model have a higher resolution in the case of [40]. Both results suggest that the QRNN model might have a higher learning potential than what was achieved in [25], and may achieve even better results for full-sized images. From 1 and 3 we find that the $28\times 28$ FRQI encoding uses 11 qubits, so the circuit will use 15 or more qubits, depending on the number of memory lanes, while the number of cells will increase to 25. The increase in the number of cells and qubits will result in a higher number of the model’s trainable parameters. However, it will also increase the model’s training capacity and potentially improve the results.
@@
The authors of [38] presented a VQDNN model, which resembles a densely connected network. They managed to heavily reduce the number of qubits required to operate the network, providing a solution that uses only ten qubits for 10-class problems. For the features extracted using the PCA, the authors experimented with amplitude and angle encoding, as well as their combination. The total number of parameters for the deepest model was 430. The depth of the model increases the model’s training capacity, and the best model achieved approximately 80% accuracy. The proposed model of the VQDNN network proved its potential for experimentation to replace the internals of the QRNN cell from [25].
@@
The use of PCA as the feature extraction method should also be considered in the further development of FRQI-QRNN models, as its impact was shown in [1] and [38]. However, the method might not be directly applicable to the problem [25] is trying to solve, i.e., the quantum ML model operating on already quantum-encoded data. Despite the fact that the PCA method can help reduce the dimensionality of feature spaces, thus reducing the model complexity and allowing for higher performance, with the same number of trainable parameters–one can imagine that the data stored in future quantum memory would be encoded and compressed using different means. The applicability of the quantum version of PCA [20, 22] or other methods such as [18] to the feature preprocessing phase should also be explored.
@@
Other works that use the amplitude encoding method for the classification of MNIST datasets are [12, 42]. They both present classifier types different from QRNN, but the authors managed to achieve high test accuracy on limited datasets. Random QNN [12] shows high robustness against noisy data. However, the image embedding is performed by a classical densely connected layer, which, similarly to PCA, offloads some part of the solution to the classical part.
@@
[42] proposed the Approximation Quantum State Preparation (AQSP) method for image encoding, which uses a simulated quantum circuit to train the image representation and has a time complexity of only $\mathcal{O}(n)$. The authors also combine the AQSP method with the proposed QCNN framework. The design of a hybrid recurrent network with convolutional and pooling input layers might help make the FRQI-QRNN architecture less dependent on the input image size and reduce the complexity of the recurrent part.
@@ V Conclusions
The new FRQI Pairs architecture, presented in [25], requires an exponentially lower number of recurrent cells compared to [1] ($n^{2}$ vs. $2^{2n}$ 3) which may lead to shorter execution times thus a lower chance of decoherence, and a higher computational efficiency during the inference phase if applied to real-world problems.
@@
Another important feat of the FRQI Pairs method is that it managed to train on the full MNIST dataset and has proven performance comparable to other methods like [22, 38]. Many other presented works struggle to capture the idea of MNIST data base, which is to create a dataset for benchmarking machine learning models and ideas based on reliable and repeatable dataset. MNIST is meant to be taken as a whole, to enable direct comparison with numerous classical models.
Although areas were left to improve compared to [1], the method has the potential to be extended by various preprocessing routines [1, 18, 20, 22] and layer/model architectures [38, 40, 42]. The approach of the other works to the problem suggests that the research from [25] should also be extended to a binary image classification case for broader comparison possibilities.